# FauForensics: Boosting Audio-Visual Deepfake Detection with Facial Action Units

### v1

@@ III-C Implicit Feature Alignment
Before modeling the temporal correlation between audio and video modalities, we attempt to utilize a novel implicit feature alignment method to learn the common information between the two modalities and reduce information redundancy for improving the generalization of audio-visual deepfake detection. Motivated by [33, 34], we design a Query-shared multimodal transformer $f_{\rm qt}(\bm{Q},\bm{K}_{\rm a},\bm{V}_{\rm a},\bm{K}_{\rm v},\bm{V}_{\rm v})=[{\bm{Z}}_{\rm aq},{\bm{Z}}_{\rm vq}]$ with weight-shared learnable queries ${\bm{Q}}\in\mathbb{R}^{T\times L}$:
Audio/visual latent features can be mapped into Key and Value (${\bm{K}\&\bm{V}}$) pairs via $1\times 1$ convolutions. Query-shared Transformer (QT) operation can be defined as follows:
@@
${\bm{Z}}_{\rm vq}={\rm softmax}({\bm{Q}}{\bm{K}_{\rm v}}^{\rm T}/\sqrt{d_{\rm v}}){\bm{V}_{\rm v}}.$ (2)
The aligned features ${\bm{Z}}_{\rm aq}\in\mathbb{R}^{T\times L}$ and ${\bm{Z}}_{\rm vq}\in\mathbb{R}^{T\times L}$ contribute to learning inter-modal feature interaction.
@@ III-D Temporal Attentional Pooler
To mine the fine-grained interaction information of both intra- and inter-modalities, we attempt to enable temporal consistency modeling by calculating frame-wise correlation within intra- and inter-modalities. The temporal consistency is expected to generalize well to various and unseen forgery methods from different modalities. To this end, we design a Temporal Attentional Pooler (TAP) that can produce a dense attentional matrix to measure the temporal consistency of intra- or inter-modalities. The dense attentional matrix ${\bm{M}}_{\rm sa}\in\mathbb{R}^{T\times T}$ can be obtained from ${\bm{Z}}_{\rm aq}$ and ${\bm{Z}}_{\rm vq}$ as follows:
@@
The above audio-visual dense attentional matrix ${\bm{M}}_{\rm av}$ can describe interaction information between two modalities. Similarly, we can obtain an audio-visual dense attentional matrix by modeling the intra-modal temporal relationship as follows:
@@
${\bm{M}}_{\rm v}=f_{\rm norm}(\sigma_{\rm v}*f_{\rm mp}({\bm{Z}}_{\rm vq},{\bm{Z}}_{\rm vq})).$ (5)
@@ III-E Multiple Forgery Detection
Next, the dense attentional matrix of inter- and intra-modalities are flattened and then fed into independent multi-layer perceptron (MLP) $f^{\rm av}_{\rm mlp}$ for forgery prediction. The inter-modality prediction ${\bm{s}_{\rm av}}$ of the audio-visual deepfake detection can be formulated as:
@@
where $f_{\rm\rho}$ indicates the flattened operation. Similarly, we can obtain visual-only and audio-only forgery predictions (i.e., ${\bm{s}_{\rm a}}$ and ${\bm{s}_{\rm v}}$) for the unimodal deepfake detection, as follows:
@@
${\bm{s}_{\rm v}}=f^{\rm v}_{\rm mlp}(f_{\rm\rho}({\bm{M}}_{\rm v})).$ (8)
Finally, we utilize general cross-entropy loss function $f_{\rm ce}(\cdot,\cdot)$ to compute the loss $\mathcal{L}_{\rm*}=f_{\rm ce}({\bm{y}},{\bm{s}})$ between Ground-Truth ${\bm{y}}$ and predictions ${\bm{s}}$. The total loss function can be obtained as follows:
@@
where $\lambda_{\rm av}$, $\lambda_{\rm a}$, and $\lambda_{\rm v}$ denotes the weight parameters for balancing these losses. The multimodal prediction score ${\bm{s}_{\rm av}}$ serves as the final decision during the inference phase.
@@ III-F Implementation Details
Specifically, our video encoder, FAU encoder, and audio encoder are obtained from the backbone network of CSN [35], ME-GraphAU [36], and Whisper [37], respectively. FAU encoder is pre-trained in the DISFA dataset [38]. The MLP module is only a single-layer fully-connected layer with 512 neurons. The number of video frames is $T=25$. The dimension of the learnable queries ${\bm{Q}}\in\mathbb{R}^{T\times L}$ is $25\times 512$.
@@ IV-A Datasets and Settings
Dataset. In this work, we evaluate the performance of the proposed method using FakeAVCeleb [39] (FakeAV for short) and LAV-DF [40] datasets for within- and cross-database evaluations.
FakeAV is a widely used multimodal forgery dataset for audio-visual deepfake detection. It is established with various visual forgery and voice cloning methods based on 500 English-speaking videos. It contains 500 real and over 20,000 fake videos. Considering the lack of an official train/test split, we use an identity-independent train/test splitting method, following [6].
LAV-DF is a novel large-scale audio-visual deepfake dataset for temporal forgery detection and localization. This dataset contains 36431 real videos and 99873 fake videos. Both FakeAV and LAV-DF consist of four audio-visual types, i.e., RealAudio-RealVideo ($RARV$), FakeAudio-RealVideo ($FARV$), RealAudio-FakeVideo ($RAFV$), and FakeAudio-FakeVideo ($FAFV$). Moreover, we utilize a famous DISFA dataset [38] to pre-train our FAU encoder. DISFA consists of 130,788 facial images with AU labels and is widely used for facial action unit recognition.
Data Preparation. To focus on mouth moving, we use center crop to obtain $160\times 160$ facial region for training our proposed model. For audio signals, we adopt log-Mel spectrograms calculated using a 25ms Hann window with the hop of 10ms and 80 Mel filter banks as acoustic input features. The resulting mel-spectrogram image is a 2D array with the size of $80\times 100$. We take 25 consecutive frames and the corresponding audio signals as the input of the model. In this work, we apply general data augmentation methods, following RealForensics [20].
TABLE I: Binary classification performance comparisons among 12 different forgery detectors in within-database evaluation.
@@
TABLE II: Four-class classification performance comparison among 4 different forgery detectors in within-database evaluation.
@@
Compared Methods. To illustrate the effectiveness of the proposed method, we compare our method with 9 advanced detectors including 4 unimodal detectors (i.e., LipForensics [25], RealForensics [20], RawGAT [27], SSLAS [11]), 3 ensemble detectors (i.e., LipForensics [25]$+$RawGAT [27], RealForensics [20]$+$RawGAT [27], and EnsembleAVDF [3]), and 4 multimodal detectors (i.e., AVD-DDL [2], AVTS [41], MRDF [6], and FGMDF [9]). AVD-DDL [2], a multimodal deepfake detector, utilizes a threshold based on modality dissonance score to predict whether the video is real or fake, which is only suitable for binary classification but not for four-class classification. Similarly, unimodal and ensemble detectors are only used for binary classification tasks.
Training Details. We implement our proposed method on the PyTorch platform and utilize AdamP [42] optimizer to update network weights. The initial learning rate, mini-batch size, and epoch are set to 1e-4, 32, and 50, respectively. Besides, we employ a poly learning rate policy with a power of 0.9 to update the learning rate.
Evaluation Metrics. In this work, we adopt two widely used metrics in the deepfake detection field, i.e., Accuracy (Acc for short) and Area Under the ROC Curve (AUC). Our proposed detector is a video-level method, so we report the experimental results of all video-level detectors based on the video clip-level ground-truth. For ensemble detectors, we average both unimodal predictions to obtain multimodal predictions.
@@ IV-B Within-database Evaluation
Most work usually considers unimodal deepfake detection as a binary classification task. However, for the multimodal deepfake detection task, simple binary classification is not enough. To fully study the types of forgery from multimodal deepfakes, we perform four-class classification (i.e., RealVideo-RealAudio, RealVideo-FakeAudio, FakeVideo-RealAudio, FakeVideo-FakeAudio) except binary classification in within-database evaluations. To fully illustrate the effectiveness of the proposed method, we compare our proposed method with previous state-of-the-art approaches in binary classification and four-class classification settings within the FakeAV and LAV-DF databases. Table I and Table II show the binary classification and four-class classification performance comparison, respectively. In both Table, the last column named “AVG” indicates the average performance between FakeAV and LAV-DF evaluations (the same as below unless indicated).
In Table I, we can clearly see that our proposed method outperforms all the detectors on the average value within FakeAV and LAV-DF. Due to the lack of modality, it is difficult for unimodal detectors to handle multimodal forgery. The ensemble-based method lacks inter-modal interaction, which limits the performance of multimodal forgery detection. By comparison, many multimodal-based detectors can detect the authenticity of videos and obtain promising performance in binary within-database evaluation. Among them, both our proposed method and FGMDF achieve outstanding average results on the two datasets for binary classification. It should be noted that FGMDF relies on 4-second video clips as the inputs, thus removing some videos shorter than 4 seconds. Fair comparisons between ours and FGMDF can be found in Fig. 3.
@@
In Table II, we show performance comparisons among 4 different forgery detectors for four-class classification on FakeAV and LAV-DF. Our proposed method also outperforms all the detectors on all metrics in more difficult four-class classification scenarios. Although FGMDF can obtain suboptimal average results in binary classification scenarios, it suffers from obvious performance degradation for four-class classification on more challenging LAV-DF database. By comparison, our proposed method can achieve optimal average performance in multiple classification scenarios. Specifically, our method outperforms the current state-of-the-art (i.e., FGMDF), achieving an average improvement of 4.41% in AUC and 12.76% in accuracy. These results illustrate that our method can accurately capture various forgeries in within-database scenarios.
@@
To further analyze experimental results, we visualize the confusion matrix of four-class classification results from the test set in four-class classification scenarios, as shown in Fig. 3. To perform a fair comparison between multiple multimodal detectors, we use the same test list as FGMDF on the LAV-DF dataset. In Fig. 3, we can see that FGMDF easily identifies $RAFV$ as $FAFV$ while our method can accurately capture the forgery of $RAFV$. These results further illustrate the superiority of the proposed approach over current advanced detectors.
IV-C Cross-database Evaluation
In the deepfake detection field, cross-database evaluation has always been the most challenging task. Cross-database evaluation means training on one database and testing on another database, which is widely used to evaluate the generalization performance of deepfake detectors against unseen forgeries and unknown post-processing operations. Generally, most deepfake detectors can obtain promising performance in within-database evaluation but usually suffer from significant performance drops in cross-database evaluation. In this section, we will focus on performance comparison in cross-database evaluation, as shown in Table III.
@@ IV-C Cross-database Evaluation
Figure 3: The confusion matrix of four multimodal forgery detectors in four-class classification evaluations. The element of the confusion matrix indicates the number of samples predicted to belong to that class. Note that four detectors use a test list of FGMDF for a fair comparison.
Table III shows the performance (AUC) comparison of the cross-database evaluation for both binary and four-class classification. The performance comparison of the two-class and four-class classification in Table 3 also contains two groups of cross-database evaluation experiments: 1) training in LAV-DF and testing in FakeAV, and 2) training in FakeAV and testing in LAV-DF. We can clearly see that our method achieves state-of-the-art results in cross-database evaluations. Compared with the prior state-of-the-art audio-visual deepfake detector, our method improves AUC by 5.74% on average (i.e., ours with 90.52% v.s. FGMDF with 84.78%) in a binary classification scenario. Similarly, in the more complex four-class classification scenario of cross-database evaluations, all the detectors meet the same performance degradation as in four-class within-database evaluations. Our method also meet this problem but obtain smaller performance degradation than other detectors, and obtain the best AUC on average. These results prove that our method benefits from the powerful feature learning ability of the FAUs-guided framework can generalize to various unseen forgeries.
Figure 4: Robustness to video compression on LAV-DF. The accuracy score (%) is reported under different compression levels.
IV-D Robustness to Video Compression
A common challenge arises from the widespread compression of videos uploaded to online platforms, as this post-processing operation degrades manipulation artifacts, compromising the robustness of forgery detection systems. In this work, we conduct a comprehensive experimental evaluation comparing our approach with two state-of-the-art methods across various video compression rates (0, 30, 32, 35, 38, 40), as shown in Fig. 4. The experimental results demonstrate that our method maintains superior robustness against video compression owing to the powerful FAUs-enhanced cross-modal modeling mechanism.

### v2

@@ III-C Implicit Feature Alignment
Before modeling the temporal correlation between audio and visual modalities, we propose a novel implicit feature alignment approach to learn the common information between the two modalities and reduce information redundancy, thereby enhancing the generalization of audio-visual deepfake detection. Motivated by [47, 48], we design a Query-shared multimodal transformer $f_{\rm t}^{\rm q}(\bm{Q},\bm{K}_{\rm a},\bm{V}_{\rm a},\bm{K}_{\rm v},\bm{V}_{\rm v})=[{\bm{Z}}_{\rm aq},{\bm{Z}}_{\rm vq}]$ with weight-shared learnable queries ${\bm{Q}}\in\mathbb{R}^{T\times L}$.
Both audio and visual latent features, ${\bm{Z}}_{\rm a}$ and ${\bm{Z}}_{\rm v}$, can be mapped into Key and Value (${\bm{K}\&\bm{V}}$) pairs using $1\times 1$ convolutions. The Query-shared Transformer (QT) operation is defined as follows:
@@
${\bm{Z}}_{\rm vq}={\rm softmax}({\bm{Q}}{\bm{K}_{\rm v}}^{\rm T}/\sqrt{d_{\rm v}}){\bm{V}_{\rm v}},$ (3)
where $\sqrt{d_{\rm v}}$ and $\sqrt{d_{\rm a}}$ serve as scaling factors. The aligned features ${\bm{Z}}_{\rm aq}\in\mathbb{R}^{T\times L}$ and ${\bm{Z}}_{\rm vq}\in\mathbb{R}^{T\times L}$ facilitate the learning of inter-modal feature interactions.
@@ III-D Temporal Attentional Pooler
To capture fine-grained interaction information across both intra- and inter-modalities, we attempt to model temporal consistency by calculating frame-wise correlations within intra- and inter-modalities. Temporal consistency is expected to generalize effectively to diverse and unseen forgery methods across different modalities. To this end, we design a Temporal Attentional Pooler that can produce a dense attentional matrix to measure the temporal consistency of intra- or inter-modalities. The dense attentional matrix ${\bm{M}}_{\rm av}\in\mathbb{R}^{T\times T}$ can be obtained from ${\bm{Z}}_{\rm aq}$ and ${\bm{Z}}_{\rm vq}$ as follows:
@@
The above audio-visual dense attentional matrix ${\bm{M}}_{\rm av}$ can describe discriminative interaction information between the two modalities, which is expected to reveal inter-modal forgery traces. Similarly, we can obtain the audio/visual dense attentional matrix, ${\bm{M}}_{\rm a}$/${\bm{M}}_{\rm v}$, by modeling the intra-modal temporal relationship as follows:
@@
${\bm{M}}_{\rm v}=f_{\rm norm}(\sigma_{\rm v}*f_{\rm mp}({\bm{Z}}_{\rm vq},{\bm{Z}}_{\rm vq})),$ (6)
where $\sigma_{\rm a}$ and $\sigma_{\rm v}$ denote scaling factors with learnable parameters for the audio and visual modalities, respectively.
@@ III-E Multiple Forgery Detection
Next, the dense attentional matrix of the inter-modality is flattened and fed into an independent multi-layer perceptron (MLP) $f^{\rm av}_{\rm mlp}$ for multimodal forgery prediction. The inter-modality prediction ${\bm{s}_{\rm av}}$ for the audio-visual deepfake detection can be formulated as:
@@
where $f_{\rm\rho}$ indicates the flattened operation. Similarly, we can obtain visual-only and audio-only forgery predictions (i.e., ${\bm{s}_{\rm a}}$ and ${\bm{s}_{\rm v}}$) for the unimodal deepfake detection. The predicted scores for the audio and visual modalities are obtained as follows:
@@
${\bm{s}_{\rm v}}=f^{\rm v}_{\rm mlp}(f_{\rm\rho}({\bm{M}}_{\rm v})),$ (9)
where $f^{\rm a}_{\rm mlp}$ and $f^{\rm v}_{\rm mlp}$ denote the audio-only classifier and visual-only classifier, respectively.
Finally, we employ the standard cross-entropy loss function $f_{\rm ce}(\cdot,\cdot)$ to compute the loss $\mathcal{L}_{\rm*}=f_{\rm ce}({\bm{y}},{\bm{s}})$ between the ground-truth labels ${\bm{y}}$ and predicted scores ${\bm{s}}$. The total loss function can be obtained as follows:
@@
where $\lambda_{\rm av}$, $\lambda_{\rm a}$, and $\lambda_{\rm v}$ denote the weight parameters used to balance the contributions of each loss.
During the inference phase, the multimodal prediction score ${\bm{s}_{\rm av}}$ is used as the final output.
@@ III-F Implementation Details
Specifically, our video encoder, FAU encoder, and audio encoder are derived from the backbone network of CSN [49], ME-GraphAU [50], and Whisper [51], respectively. The FAU encoder is pre-trained on the DISFA dataset [52], which provides fine-grained FAU annotations essential for learning subtle facial motion patterns. The MLP module is only a single-layer fully-connected layer with 512 neurons. In our framework, each input video sample consists of $T=25$ frames. The dimension of the learnable queries ${\bm{Q}}\in\mathbb{R}^{T\times L}$ used in our Query-shared Transformer is $25\times 512$. The trade-off parameters $\lambda_{\rm a}$, $\lambda_{\rm v}$, and $\lambda_{\rm av}$, which balance the contributions of the audio, visual, and audio-visual losses, are empirically set to $0.1$, $0.1$, and $0.8$, respectively.
@@ IV Experiments
In this section, we evaluate the performance of our proposed method on two publicly available audio-visual deepfake datasets (i.e., LAV-DF [53] and FakeAVCeleb [54]) for within- and cross-database evaluations.
@@ IV-A Datasets and Settings
Dataset. FakeAVCeleb (FakeAV for short) is a widely used multimodal forgery dataset for audio-visual deepfake detection. It is constructed using a combination of visual manipulation techniques and voice cloning methods, based on 500 English-speaking video samples. The dataset comprises 500 real videos and over 20,000 forged ones. Due to the absence of an official train/test split, we adopt an identity-independent train/test splitting strategy, following the protocol described in [7].
LAV-DF is a newly introduced large-scale audio-visual deepfake dataset established for temporal forgery detection and localization. It contains 36,431 real videos and 99,873 fake videos, and provides an official train/test split for comprehensive performance evaluation of detectors.
Both the FakeAV and LAV-DF datasets include four types of audio-visual combinations: RealAudio-RealVideo ($RARV$), FakeAudio-RealVideo ($FARV$), RealAudio-FakeVideo ($RAFV$), and FakeAudio-FakeVideo ($FAFV$). Representative samples from these four categories are illustrated in Fig. 3. In addition, we leverage the well-known DISFA dataset [52] to pre-train our FAU encoder. DISFA consists of 130,788 facial images annotated with action unit (AU) labels and is widely adopted for facial action unit recognition tasks.
Figure 3: Samples from four categories of audio-visual signals. Forged samples are highlighted with a red box.
Data Preparation. To focus on mouth movements, we apply a center crop to extract a $160\times 160$ facial region for training our proposed model. For audio processing, we compute log-Mel spectrograms using a 25-ms Hann window with a hop of 10 ms and 80 Mel filter banks as acoustic input features. The resulting mel-spectrogram is a 2D array with the size of $80\times 100$. We take 25 consecutive video frames along with the corresponding audio signals as the input of the model. In addition, we employ general data augmentation methods, following RealForensics [29].
TABLE I: Binary classification performance comparisons among 12 different forgery detectors on FakeAV and LAV-DF under within-database evaluations.
@@
TABLE II: Four-class classification performance comparison among 4 different forgery detectors on FakeAV and LAV-DF under within-database evaluations.
@@
Compared Methods. To demonstrate the effectiveness of the proposed method, we compare our method with 9 advanced deepfake detectors including 4 unimodal detectors (i.e., LipForensics [36], RealForensics [29], RawGAT [55], SSLAS [56]), 3 ensemble detectors (i.e., LipForensics [36]$+$RawGAT [55], RealForensics [29]$+$RawGAT [55], and EnsembleAVDF [3]), and 4 multimodal detectors (i.e., AVD-DDL [2], AVTS [57], MRDF [7], and FGMDF [10]). It should be noted that AVD-DDL [2], a multimodal deepfake detector, predicts authenticity based on a modality dissonance score threshold, making it suitable only for binary classification tasks and unsuitable for four-class classification. Similarly, the unimodal and ensemble detectors are also limited to binary classification and can not be directly applied to four-class classification scenarios.
Training Details. We implement our proposed method on the PyTorch platform and utilize the AdamP [58] optimizer to update the network weights. The initial learning rate, mini-batch size, and number of training epochs are set to 1e-4, 32, and 50, respectively. Besides, we employ a poly learning rate policy with a power of 0.9 to adjust the learning rate during training.
Evaluation Metrics. In this work, we adopt two widely used evaluation metrics in the deepfake detection field, i.e., Accuracy (Acc for short) and Area Under the ROC Curve (AUC). Since our proposed method operates at the video level, we report experimental results for all video-level detectors based on the video clip-level ground truth. For ensemble detectors, we obtain multimodal predictions by averaging the outputs of the corresponding unimodal detectors.
@@ IV-B Within-database Evaluation
Most existing work usually treats unimodal deepfake detection as a binary classification task, where the objective is to distinguish between authentic and forged content. However, since audio and visual modalities can be manipulated either independently or jointly in multimodal deepfakes, simple binary classification is insufficient. This is because binary classification fails to provide insights into the specific category of multimodal forgery. As a result, it restricts the detector’s interpretability and reduces its usefulness in forensic analysis, where understanding the exact type of manipulation is crucial for informed decision-making and further investigation. To comprehensively analyze the types of forgery in multimodal deepfakes, we conduct both binary and four-class classification in within-database and cross-database evaluation. The four-class classification setting includes the categories: RealVideo–RealAudio, RealVideo–FakeAudio, FakeVideo–RealAudio, and FakeVideo–FakeAudio. Table I and Table II report the binary and four-class classification results on the FakeAV and LAV-DF datasets, respectively. In both tables, the last column labeled “AVG” represents the average performance across FakeAV and LAV-DF evaluations (the same as below unless indicated).
As shown in Table I, we compare our proposed method with four unimodal detectors, three ensemble-based detectors, and four multimodal detectors in within-database evaluations. We can see that our proposed method outperforms all other detectors in terms of average performance across the FakeAV and LAV-DF datasets. Unimodal detectors struggle with multimodal forgeries due to the absence of multimodal information, making it difficult to capture all the artifacts from both audio and visual streams. Although ensemble-based detectors aggregate predictions from unimodal models, they lack explicit modeling of inter-modal interactions, which hampers their ability to detect complex cross-modal manipulations effectively. In contrast, multimodal detectors are more capable of capturing cross-modal inconsistencies, leading to stronger performance in binary within-database evaluations. Notably, both our method and FGMDF achieve top average results for binary classification across both datasets. However, it is worth noting that FGMDF processes 4-second video clips, which leads to the exclusion of shorter videos during evaluation. A fairer comparison between our approach and FGMDF, accounting for these differences, is presented in Fig. 4.
@@
Table II presents performance comparisons among 4 different forgery detectors for the four-class classification task on the FakeAV and LAV-DF datasets. Our proposed method also outperforms all the deepfake detectors across all evaluation metrics in four-class classification scenarios. Although FGMDF achieves competitive results in binary classification scenarios, it suffers from obvious performance degradation for four-class classification, particularly on the LAV-DF dataset. In contrast, our method maintains robust performance and achieves the best overall results in multiple classification scenarios. Specifically, compared to the current state-of-the-art (i.e., FGMDF), our approach yields an average improvement of 4.41% in AUC and 12.76% in accuracy. These results demonstrate the effectiveness of our method in accurately detecting various types of forgeries under within-database evaluation settings.
@@
Figure 4: The confusion matrix of four multimodal forgery detectors in four-class classification evaluations. Each element in the matrix represents the number of samples predicted for the corresponding class. For a fair comparison, all four detectors are evaluated using the same test list employed by FGMDF.
To gain deeper insights into the experimental results, we visualize the confusion matrix of four-class classification results on the test set under the within-database evaluation setting, as shown in Fig. 4. To perform a fair comparison between multiple multimodal detectors, we use the same test list as FGMDF on the LAV-DF dataset. As illustrated in Fig. 4, we can see that FGMDF easily misclassifies $RAFV$ samples as $FAFV$, indicating a limitation in distinguishing between visual-only forgeries and fully forged audio-visual content. In contrast, our method achieves a much higher accuracy in identifying $RAFV$ instances. These observations further highlight the superiority of our approach over existing state-of-the-art detectors in capturing complex multimodal forgery patterns.
IV-C Cross-database Evaluation
Cross-database evaluation has been widely regarded as one of the most challenging tasks in the field of deepfake detection. Unlike within-database evaluation, which assesses the performance of detectors on the same dataset used for both training and testing, cross-database evaluation involves training on one dataset and evaluating on another unseen dataset. This setting is recognized as a reliable benchmark for evaluating the generalization ability of deepfake detectors, especially their robustness to previously unseen manipulation techniques and post-processing operations. Generally, most deepfake detectors can achieve promising results under within-database conditions but tend to exhibit substantial performance degradation when applied to new datasets. This performance drop indicates that deepfake detectors are likely overfitting to dataset-specific artifacts rather than learning generalizable forgery-relevant features. In this section, we provide a detailed comparison of different audio-visual deepfake detectors under cross-database evaluation, as summarized in Table III. Our analysis focuses on the models’ ability to maintain high detection performance when faced with real-world variability, including diverse forgery patterns and different dataset domains.
Table III presents the performance (AUC) comparison of the cross-database evaluation under both binary and four-class classification settings. The experiments include two cross-database scenarios: 1) training on LAV-DF and testing on FakeAV, and 2) training on FakeAV and testing on LAV-DF. As shown in the table III, our proposed method achieves state-of-the-art performance across both classification settings. Specifically, in the binary classification scenario, our method outperforms the previous SOTA audio-visual deepfake detector (i.e., FGMDF) by an average AUC improvement of 5.74% (90.52% vs. 84.78%). In the more challenging four-class classification scenario, all detectors suffer from an obvious performance drop due to the increased task complexity, similar to observations in within-database evaluations. However, our method exhibits a smaller performance degradation compared to other detectors and still achieves the highest average AUC across evaluations. These results demonstrate that our FAUs-guided framework not only captures subtle multimodal inconsistencies effectively but also enhances cross-dataset generalizability. This robustness against unseen forgeries and dataset domain further highlights the generalizability and reliability of our approach in real-world applications.
Figure 5: (a) Examples of facial images under different perturbations. (b) Examples of video compression perturbations at different intensity levels. Higher intensity levels indicate stronger perturbations.
Figure 6: Robustness to unseen perturbation on FakeAV. AUC scores (%) are reported across five intensity levels for six types of perturbations: video compression, JPEG compression, color contrast, Gaussian blur, Gaussian noise, and color saturation.
IV-D Robustness to Perturbations
A common challenge arises from widespread post-processing perturbations, such as image/video compression, noise, etc. These perturbation operations can degrade manipulation artifacts, thereby compromising the robustness and reliability of forgery detection systems. In this study, we consider the six perturbation methods, including Video Compression, JPEG Compression, Color Contrast, Gaussian Blur, Gaussian Noise, and Color Saturation, as presented in Fig. 5(a). Following  [59], we systematically apply each perturbation to the original video at five different intensity levels. Fig. 5(b) presents examples of video compression at different intensity levels. Note that Intensity level 0 corresponds to the original video without any perturbations. To evaluate robustness against perturbations, we conduct a comprehensive experimental evaluation comparing our approach with two prior SOTA methods across various types of perturbations, as shown in Fig. 6. The experimental results clearly demonstrate that our method achieves excellent robustness against all perturbation types and intensity levels. The superior robustness of our method stems from the FAU-enhanced cross-modal modeling mechanism, which facilitates effective feature learning even under severe post-processing distortions.